OpenStudy (anonymous):

Jack usually mows his lawn in 7hrs. Marilyn can mow the same yard in 5hrs. How much time would it take for them to mow the lawn together? They could mow the lawn in ? hrs if they worked together

OpenStudy (anonymous):

6hrs??? maybe

OpenStudy (anonymous):

Is that a guess?

OpenStudy (anonymous):

i need a good answer

OpenStudy (anonymous):

2 hours 55 min

OpenStudy (anonymous):

thx u sure

OpenStudy (anonymous):

I used algebra, so I'm pretty sure that's right

OpenStudy (anonymous):

na actually try this jack is (1lot/7hrs) where marilyn is (1lot/5hrs) adding these together gives you .343 lots per hour so 1 lot/.343 lots per hours gives you 2.96 hours

OpenStudy (anonymous):

ok how do i make that a integer or improper fraction or mixed number

OpenStudy (anonymous):

it wouldn't make sense to take 6 hours if tweo people are working on the same area

OpenStudy (anonymous):

oh nvm it was already answered... my bad :/

OpenStudy (anonymous):

Yeah too bad your 2.96 isn't in terms of hours and minutes

OpenStudy (anonymous):

obviously

OpenStudy (anonymous):

You have to convert it

OpenStudy (anonymous):

how the heck do i do that?

OpenStudy (anonymous):

2+ .96(60)

OpenStudy (anonymous):

Jack mows 1/7 lawns in an hour and Marilyn mows 1/5. Combined rate is 1/7+1/5 lawns per hour. t is the time it takes to mow lawn together, so we have (1/7+1/5)t=1 Giving t=35/12 hours=2 Hr, 55 min....Nicely done folks.

OpenStudy (anonymous):

Yeah, I'm going with 2hrs 55 min

OpenStudy (anonymous):

i have to many helping me here hero can help me

OpenStudy (anonymous):

how to convert

OpenStudy (anonymous):

Actually, the method I used only requires converting fractions into minutes, which is much easier in my opinion

OpenStudy (anonymous):

Go Blue posted the steps College girl

OpenStudy (anonymous):

so 2hr and 55min is converted already

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Unless they want it all in terms of hours....Then it would be 2 and 11/12 hours

OpenStudy (anonymous):

integer, improper fraction or mixed number

OpenStudy (anonymous):

2 11/12 then

OpenStudy (anonymous):

\[2 {{11} \over {12}}\]

OpenStudy (anonymous):

Yes